Integrate log(1/x – 1)
upper limit =1
lower limit=0

Grade:12

2 Answers

Arun
6 years ago
Dear Suraj
 
log (1/(x-1))
 = Log 1 - log (x-1)
 = 0 - log (x-1)
 
Now integrate it by byparts
And then put limit
 
Regards
Arun (askIITians forum expert)
Rohit Sharma
6 years ago
log(1/x-1)= log1-log(x-1)= 0-log(x-1)= -log(x-1)
Then either integrate it by parts or you can substitute x-1 =t and then change the limits accordingly.